# A Survey of LLM-based Automated Program Repair: Taxonomies, Design Paradigms, and Applications

### v1

@@ 2.5. Context-Enriched Fine-Tuning
InferFix couples Meta’s Infer static analyzer with retrieval-augmented fine-tuning (Jin et al., 2023). For every alert, it retrieves analogous historical patches from a large corpus, concatenates the alert description, the buggy code, and the retrieved examples, and fine-tunes a 1.3B parameter transformer to emit a patch that clears the analyzer warning. The additional context leads to 59 unique repairs that baseline LLMs miss and reduces false positives by filtering candidates that still trigger Infer.
@@
NTR (Neural Template Repair) formalizes template-based repair as a conditional text-to-text task where the selected template identifier is injected as a control token (Huang et al., 2024b). A lightweight classifier first ranks candidate templates, then a billion-parameter decoder is fine-tuned to translate the buggy code and template hint into a concrete patch. An OtherTemplate token lets the model learn fixes beyond the known catalogs. NTR with StarCoder repairs 128 bugs on Defects4J-v2 and 129 on HumanEval-Java, surpassing the best prior work while remaining template-aware.
@@
2.6. Discussion and Outlook
Research on fine-tuning for APR reveals that full supervised adaptation yields the strongest patch accuracy, as it embeds rich code knowledge. However, this benefit comes at a high GPU cost and the potential erosion of broader coding abilities. Parameter-efficient tuning recovers much of that accuracy by updating small adapter modules and low-rank matrices. Still, its effectiveness hinges on sensible adapter design and diverse, well-curated training data. Reinforcement learning introduces execution rewards that guide models toward functional correctness, yet sparse signals and repeated test runs extend training time. Supplying static analysis, execution traces, or retrieved historical patches broadens the LLM’s perspective. It helps resolve multi-file or semantic defects, though collecting and encoding such context is expensive and can exceed current input limits. The most promising direction is to combine these strengths: coupling lightweight adapters with selective retrieval of high-value context, applying reinforcement learning only when tests provide clear feedback, and adopting staged curricula with locate-edit-verify workflows to shorten feedback loops. Privacy-preserving federated or continual updates can keep models current without exposing proprietary code. Advances across model architecture, training signals, and deployment practices will be needed before fine-tuned repair systems can deliver reliable, low-latency fixes in everyday continuous integration environments.
@@ 3. Prompting Approaches
Prompt-based APR treats the LLM as a frozen component that returns a complete patch after a single query, while every step of retrieval, example curation, or validation runs outside the model. Four practical prompt styles emerge, distinguished by the amount and source of extra context they add. Zero-shot prompting supplies only the buggy code and a brief fix instruction, counting on the LLM’s prior knowledge to generate a patch (Xia and Zhang, 2022; Prenner et al., 2022; Tian et al., 2023). Few-shot prompting precedes the defect with a handful of similar bug-fix pairs, enabling the model to imitate their edit pattern (Xia et al., 2023; Gao et al., 2023; Ahmed and Devanbu, 2023; Nashid et al., 2023). Retrieval-augmented prompting attaches code fragments, templates, or knowledge-graph facts fetched from external resources, grounding the patch in concrete project or domain information (Chen et al., 2024; Liu et al., 2024b; Ouyang et al., 2025). Analysis-augmented prompting injects diagnostics derived from static or dynamic analysis, such as failing-test logs or semantics-aware slices, to guide the model toward the root cause (Xu et al., 2024; Nong et al., 2024; Haque et al., 2025b). These approaches share the same single-turn interaction but trade off preparation effort, token budget, and repair accuracy in different ways.
@@ 3.1. Zero-shot Prompting
Zero-shot prompting keeps a pre-trained model frozen and asks it to generate a patch in a single turn by providing only the buggy code and a terse instruction such as “fix the bug”. Because the approach omits fine-tuning, retrieval, and multi-turn dialogue, deployment is rapid and incurs no additional training cost. AlphaRepair (Xia and Zhang, 2022) was the first study to demonstrate that LLMs can outperform state-of-the-art APR techniques with zero-shot. Empirical studies such as Prenner et al. (Prenner et al., 2022) and Fan et al. (Fan et al., 2023a) on Codex and Tian et al. (Tian et al., 2023) on ChatGPT, show that careful wording plus a sufficiently large model can already match or sometimes surpass several traditional repair pipelines. At the same time, they expose strong prompt sensitivity, but their robustness beyond single-hunk functions remains unclear.
AlphaRepair was the first LLM-based APR approach that performs repair in a zero-shot setting by exploiting the masked-language-modeling capability of CodeBERT without any task-specific fine-tuning (Xia and Zhang, 2022). AlphaRepair masks the tokens of the buggy lines and leverages CodeBERT to infill the correct code, thereby predicting the fixed codes directly instead of learning edit scripts. On Defects4J-v1.2/2.0, AlphaRepair repairs up to 3.3 times more bugs than the best baselines, and it also outperforms other neural and heuristic tools on QuixBugs, demonstrating that a frozen pre-trained LLM can surpass state-of-the-art APR techniques without further training.
Fan et al. conducted a large-scale evaluation of zero-shot Codex on competitive-programming tasks and then studied how APR can salvage its failures (Fan et al., 2023a). Their prompt contained only the LeetCode problem description and the target function’s signature; no exemplar fixes were provided. By generating 50 solutions per task, Codex solved 46 of the 113 tasks in their new LMDefects benchmark, while subsequent APR passes (TBar, Recoder, Codex-edit) raised the total by a further 9 tasks, showing that LLM with naive zero-shot prompting already covers a substantial fraction of real bugs and that lightweight repair can extend this reach.
Prenner et al. evaluated Codex (12B) (Chen et al., 2021) on all 40 QuixBugs (Lin et al., 2017) bugs in Python and Java (Prenner et al., 2022). They tried five one-shot prompt variants that differed only in the static text preceding the buggy function. With temperature 0 and max_tokens=1024, Codex generated ten candidates per bug; a patch was accepted if QuixBugs’ reference tests all passed. The best prompt repaired 23 Python and 14 Java bugs, outperforming three specialized neural APR systems despite Codex never having seen any examples. The study also showed that adding a simple “### bug is here” hint sometimes hurts accuracy, illustrating how small prompt edits can swing outcomes.
@@ 3.2. Few-shot Prompting
Xia et al. performed the first large-scale study that applies nine modern LLMs to APR without fine-tuning (Xia et al., 2023). They evaluated three prompt settings—complete function generation, code infilling with prefix and suffix, and single-line edits, over 1,094 bugs drawn from Defects4J 1.2/2.0 (Java) (Just et al., 2014), QuixBugs (Java & Python) (Lin et al., 2017), and ManyBugs (C) (Le Goues et al., 2015). For few-shot generation, they prefixed two exemplar bug-fix pairs ahead of the target defect and requested up to 200 samples per bug. Under this setting, Codex repaired sixty-three single-function Java bugs on Defects4J, significantly better than the previous best APR tool. The study further reveals a clear scaling trend (larger models fix more bugs), highlights the value of including suffix context during infilling, and demonstrates that simple entropy scores enable frozen LLMs to rank patches effectively. These findings suggest that even a minimal number of few-shot prompts, when paired with sufficient sampling, can rival specialized learning-based systems. That prompt design should balance exemplar relevance, context length, and sample budget.
Gao et al. designed more than two hundred Codex prompts for Java repair (Gao et al., 2023). Each prompt began with exactly four bug-fix pairs that were retrieved by lexical similarity and copied verbatim, in the same order, followed by the new buggy snippet and the instruction “produce a fixed version”. Under this setting, exact-match accuracy more than doubled relative to random exemplars and tripled relative to zero-shot. The authors concluded that prompts should prioritize highly similar examples, keep their order stable, and stop adding demonstrations once the prompt nears the context limit; otherwise, gains vanish when no close fixes exist.
Ahmed et al. prefixed eight bug-fix demonstrations, each followed by its original commit message that briefly explained the change (Ahmed and Devanbu, 2023). After these exemplars, they placed the buggy JavaScript function and asked the frozen code-davinci-002 model to return a corrected version. They sampled fifty completions and selected the patch that appeared most often. The commit messages provided lightweight reasoning traces that helped the model generalize, and majority voting raised repair accuracy by roughly ten percentage points. This approach increases token usage and assumes that informative commit logs are available.
@@ 3.3. Retrieval-Augmented Generation Enhanced Prompting
Retrieval-augmented generation (RAG) enriches a single-turn repair prompt with external artifacts fetched on demand. By retrieving repository structure information (Chen et al., 2024), security-fix templates (Liu et al., 2024b), or domain-specific API knowledge encoded in a graph (Ouyang et al., 2025), a frozen language model can ground its patch in concrete evidence instead of relying solely on parametric memory. Compared with plain zero- or few-shot prompts, RAG improves factual grounding and supplies missing context while deployment remains as light as a one-shot generation.
RLCE targets repository‑level bugs that span multiple files. A static parser builds a lightweight project map and then retrieves only those out‑of‑file definitions and usages that interact with the faulty code; the selected fragments are appended to the prompt that asks GPT‑3.5 or GPT‑4 to “produce the fixed version” (Chen et al., 2024). This single‑shot context expansion raises GPT‑3.5’s top-1 fix rate on their built 124‑bug RepoBugs suite from 22.6% to 56.0% and lifts GPT‑4 from 41.1% to 75.8%, more than doubling accuracy without extra inference rounds or model updates.
T-RAP focuses on vulnerability repair. It first abstracts historical security fixes into edit-script templates (Liu et al., 2024b), in a way similar to traditional repair techniques such as HDRepair (Le et al., 2016). Given a new bug, the system retrieves the most relevant template and its concrete instance from a patch corpus, appends both to the buggy code, and asks a frozen code LLM (demonstrated with CodeT5 and GPT family models) to instantiate the fix in one pass. Experimental results show that T-RAP repairs 56.8% more vulnerabilities than the previous state-of-the-art VulRepair (Fu et al., 2022) under identical evaluation settings. The unique ingredient is template guidance: retrieval supplies a high-level edit pattern that narrows the search space, enabling the LLM to generate correct patches for complex security flaws without any iterative feedback.
DSrepair retrieves structured API knowledge from a purpose‑built Data‑Science Knowledge Graph and injects the retrieved triples plus the localized buggy snippet into a single repair prompt for GPT‑3.5, GPT‑4, and two open-source code LLMs, DeepSeek-Coder (Guo et al., 2024) and Codestral (AI, 2024), without any fine‑tuning (Ouyang et al., 2025). On the DS‑1000 benchmark (Lai et al., 2023), it fixes 44.4%, 14.2%, 20.6%, and 32.1% more data‑science bugs than the best prior baseline while cutting prompt tokens by up to 34%, showing that knowledge‑graph evidence can outperform unstructured text retrieval in one‑shot repair.
@@ 3.4. Analysis-Augmented Generation Enhanced Prompting
Analysis-augmented generation (AAG) feeds a frozen language model with artifacts from program analysis, such as compiler diagnostics, fault localization scores, data flow constraints, failing test logs, or execution traces. These artifacts typically appear in three forms: failing-test evidence that exposes runtime failures, semantics-aware slices that prune unrelated code, and concise execution traces that embed dynamic behavior. Recent systems illustrate each form in turn: D4C (Xu et al., 2024) injects failing-test evidence, Appatch (Nong et al., 2024) applies semantics-aware slices, and TracePrompt (Haque et al., 2025b) contributes concise execution traces. Each patch is grounded in concrete evidence and is generated in a single pass. Compared to plain zero-shot prompts, AAG consistently improves repair accuracy and keeps edits focused on fault-relevant code, while ensuring that extra context remains within the model window.
D4C aligns the prompt with decoder-only models’ next-token objective and lets GPT-4 complete the whole function while embedding the failing test case and its output (Xu et al., 2024). Without any fault-localization hints, D4C fixes 180 of 437 single-function bugs in Defects4J, 10% more than the previous state of the art that assumed perfect localization, with only ten samples per bug. The approach relies on dynamic evidence, so its gains diminish when tests give little diagnostic detail.
Appatch applies semantics-aware scoping and adaptive exemplar prompting to vulnerability repair (Nong et al., 2024). It slices the program to include only code relevant to the vulnerability, mines an exemplar database, and selects examples on the fly to guide patch generation. In evaluations on 306 real vulnerabilities, it raised F1 from 44.9% to 57.2% across four LLMs, outpacing baseline prompting by up to 98.9% and surpassing state-of-the-art static approaches by as much as 65.4%. Appatch also fixed seven of eleven zero-day vulnerabilities that none of the baselines could handle. Strengths include reduced context, tailored exemplars, and ensemble validation across models; limitations are the need for a known vulnerability location and extra token overhead from exemplar insertion.
@@
3.5. Discussion and Outlook
Prompt-based repair has evolved from succinct zero-shot instructions to prompts enriched with exemplar fixes, retrieved project fragments, and static or dynamic analysis outputs. Larger frozen language models consistently repair more bugs without requiring parameter updates, and they become even more effective when supplied with focused auxiliary information, such as highly similar bug-fix pairs, carefully selected repository or API snippets, or compact failing-test logs and execution traces. Conversely, verbose or loosely related context can distract the model and harm accuracy, underscoring the need to balance relevance against prompt length. Most current pipelines still rely on manual wording or heuristic retrieval, but evidence from CEDAR (Nashid et al., 2023) shows that automatic exemplar selection and prompt compression are already practical. Continued research on adaptive artifact gathering, scalable summarization, and systematic evaluation will be essential for making prompt-centric repair a dependable component of continuous-integration workflows and developer tooling.

### v2

@@ 4.5. Context-Enriched Fine-Tuning
InferFix couples Meta’s Infer static analyzer with retrieval-augmented fine-tuning (Jin et al., 2023). For every alert, it retrieves analogous historical patches from a large corpus, concatenates the alert description, the buggy code, and the retrieved examples, and fine-tunes Codex-12B to generate patches that clear the analyzer warning. The additional context leads to 59 unique repairs that baseline LLMs miss and reduces false positives by filtering candidates that still trigger Infer.
@@
NTR (Neural Template Repair) formalizes template-based repair as a conditional text-to-text task where the selected template identifier is injected as a control token (Huang et al., 2024b). A lightweight classifier first ranks candidate templates, then a billion-parameter decoder is fine-tuned to translate the buggy code and template hint into a concrete patch. An OtherTemplate token lets the model learn fixes beyond the known catalogs. NTR with CodeLlama-70B repairs 139 bugs on Defects4J v2 and 136 on HumanEval-Java, surpassing the best prior work while remaining template-aware.
@@
Overview Research on fine-tuning for software repair shows a clear tradeoff between accuracy, cost, and generality. Full supervised adaptation still delivers the strongest patch accuracy, but is expensive and can erode broader coding abilities. Parameter-efficient tuning and distillation recover much of this accuracy at far lower training and inference cost, yet depend on careful adapter or student design and diverse, high-quality bug-fix data. Reinforcement- and preference-based objectives push models toward execution-level correctness but remain sensitive to reward design and test availability, while context-enriched training that incorporates analysis results, traces, or retrieved patches improves multi-file and semantic defects at the price of heavy pre-processing and context management. A promising direction is to combine these strengths in cost-aware curricula and privacy-preserving continual updates, so that fine-tuned repair models can offer reliable, low-latency fixes that fit continuous integration constraints.
@@ 5. Prompting Approaches
Prompt-based software repair treats the LLM as a frozen component that returns a complete patch after a single query, while every step of retrieval, example curation, or validation runs outside the model. Four practical prompt styles emerge, distinguished by the amount and source of extra context they add. Zero-shot prompting supplies only the buggy code and a brief fix instruction, counting on the LLM’s prior knowledge to generate a patch (Xia and Zhang, 2022; Prenner et al., 2022; Tian et al., 2023). Few-shot prompting precedes the defect with a handful of similar bug-fix pairs, enabling the model to imitate their edit pattern (Xia et al., 2023; Gao et al., 2023; Ahmed and Devanbu, 2023; Nashid et al., 2023). Retrieval-augmented prompting attaches code fragments, templates, or knowledge-graph facts fetched from external resources, grounding the patch in concrete project or domain information (Chen et al., 2024; Ehsani et al., 2025; Ouyang et al., 2025). Analysis-augmented prompting injects diagnostics derived from static or dynamic analysis, such as failing-test logs or semantics-aware slices, to guide the model toward the root cause (Xu et al., 2025a; Nong et al., 2025; Haque et al., 2025b). These approaches share the same single-turn interaction but trade off preparation effort, token budget, and repair accuracy in different ways.
@@ 5.1. Zero-shot Prompting
Zero-shot prompting keeps a pre-trained model frozen and asks it to generate a patch in a single turn, given only the buggy code and a terse instruction such as “fix the bug”. Because the approach omits fine-tuning, retrieval, and multi-turn dialogue, deployment is rapid and incurs no additional training cost. AlphaRepair (Xia and Zhang, 2022) was the first study to demonstrate that LLMs can outperform state-of-the-art repair techniques with zero-shot. Empirical studies such as Prenner et al. (Prenner et al., 2022) and Fan et al. (Fan et al., 2023a) on Codex and Tian et al. (Tian et al., 2023) on ChatGPT show that careful wording plus a sufficiently large model can already match or sometimes surpass several traditional repair pipelines. At the same time, they expose strong prompt sensitivity, but their robustness beyond single-hunk functions remains unclear.
AlphaRepair (Xia and Zhang, 2022) masks the tokens of the buggy lines and leverages CodeBERT to infill the correct code, thereby predicting the fixed codes directly instead of learning edit scripts. On Defects4J-v1.2/2.0, AlphaRepair repairs up to 3.3 times more bugs than the best baselines, and it also outperforms other neural and heuristic tools on QuixBugs, demonstrating that a frozen pre-trained LLM can surpass state-of-the-art repair techniques without further training.
Prenner et al. evaluated Codex (12B) (Chen et al., 2021) on all 40 QuixBugs (Lin et al., 2017) bugs in Python and Java (Prenner et al., 2022). They tried five one-shot prompt variants that differed only in the static text preceding the buggy function. With temperature 0 and max_tokens=1024, Codex generated ten candidates per bug; a patch was accepted if QuixBugs’ reference tests all passed. The best prompt repaired 23 Python and 14 Java bugs, outperforming three specialized neural repair systems despite Codex never having seen any examples. The study also showed that adding a simple “### bug is here” hint sometimes hurts accuracy, illustrating how small prompt edits can swing outcomes.
Fan et al. conducted a large-scale evaluation of zero-shot Codex on competitive-programming tasks and then studied how software repair can salvage its failures (Fan et al., 2023a). Their prompt contained only the LeetCode problem description and the target function’s signature; no exemplar fixes were provided. By generating 50 solutions per task, Codex solved 46 of the 113 tasks on their new LMDefects benchmark, showing that LLM with naive zero-shot prompting already covers a substantial fraction of real bugs and that lightweight repair can extend this reach.
@@ 5.2. Few-shot Prompting
Xia et al. performed the first large-scale study that applies nine modern LLMs to repair without fine-tuning (Xia et al., 2023). They evaluated three prompt settings—complete function generation, code infilling with prefix and suffix, and single-line edits, over 1,094 bugs drawn from Defects4J 1.2/2.0 (Java) (Just et al., 2014), QuixBugs (Java & Python) (Lin et al., 2017), and ManyBugs (C) (Le Goues et al., 2015). For few-shot generation, they prefixed two exemplar bug-fix pairs ahead of the target defect and requested up to 200 samples per bug. Under this setting, Codex repaired sixty-three single-function Java bugs on Defects4J, significantly better than the previous best repair tool. The study further reveals a clear scaling trend (larger models fix more bugs), highlights the value of including suffix context during infilling, and demonstrates that simple entropy scores enable frozen LLMs to rank patches effectively. These findings suggest that even a minimal number of few-shot prompts, when paired with sufficient sampling, can rival specialized learning-based systems. That prompt design should balance exemplar relevance, context length, and sample budget.
Gao et al. designed more than two hundred Codex prompts for Java repair (Gao et al., 2023). Each prompt began with exactly four bug-fix pairs that were retrieved by lexical similarity and copied verbatim, in the same order, followed by the new buggy snippet and the instruction for producing a fixed version. Under this setting, exact-match accuracy more than doubled relative to random exemplars and tripled relative to zero-shot. The authors concluded that prompts should prioritize highly similar examples, keep their order stable, and stop adding demonstrations once the prompt nears the context limit; otherwise, gains vanish when no close fixes exist.
Ahmed et al. prefixed eight bug-fix demonstrations, each followed by its original commit message that briefly explained the change (Ahmed and Devanbu, 2023). After these exemplars, they placed the buggy JavaScript function and asked the frozen code-davinci-002 model to return a corrected version. They sampled 30 completions and selected the patch that appeared most often. The commit messages provided lightweight reasoning traces that helped the model generalize, and majority voting raised repair accuracy by roughly ten percentage points. This approach increases token usage and assumes that informative commit logs are available.
@@ 5.3. Retrieval-Augmented Generation Enhanced Prompting
Retrieval-augmented generation (RAG) enriches a single-turn repair prompt with external artifacts fetched on demand. By retrieving repository structure information (Chen et al., 2024), layered knowledge injection (Ehsani et al., 2025), or domain-specific API knowledge encoded in a graph (Ouyang et al., 2025), a frozen language model can ground its patch in concrete evidence instead of relying solely on parametric memory. Compared with plain zero- or few-shot prompts, RAG improves factual grounding and supplies missing context while deployment remains as light as a one-shot generation.
RLCE targets repository-level bugs that span multiple files. A static parser builds a lightweight project map and then retrieves only those out-of-file definitions and usages that interact with the faulty code; the selected fragments are appended to the prompt that asks GPT-3.5 or GPT-4 to produce the fixed version (Chen et al., 2024). This single-shot context expansion raises GPT-3.5’s top-1 fix rate on their built 124-bug RepoBugs suite from 22.6% to 56.0% and lifts GPT-4 from 41.1% to 81.5%, more than doubling accuracy without extra inference rounds or model updates.
Ehsani et al. propose a layered knowledge injection framework for LLM-based software repair that incrementally augments prompts with Bug Knowledge, Repository Knowledge, and Project Knowledge automatically extracted from the buggy function, co-changing files, Git history, documentation, and previously resolved issues (Ehsani et al., 2025). On the BugsInPy (Widyasari et al., 2020) benchmark of 314 Python bugs, this framework with Llama 3.3 raises the fix rate to 79% by correctly repairing 250 bugs, a 23% improvement over prior work, with 65% fixed using only the Bug Knowledge Layer, a further 9% gain from adding the Repository Knowledge Layer, and an additional 5% gain from the Project Knowledge Layer. They find that retrieved context should be injected hierarchically and in a bug-type-aware way, rather than being concatenated all at once, because different bug types benefit from different kinds and depths of retrieved repository and project knowledge for effective repair.
DSrepair retrieves structured API knowledge from a purpose-built Data-Science Knowledge Graph and injects the retrieved triples plus the localized buggy snippet into a single repair prompt for GPT-3.5, GPT-4, and two open-source code LLMs, DeepSeek-Coder (Guo et al., 2024) and Codestral (AI, 2024), without any fine-tuning (Ouyang et al., 2025). On the DS-1000 benchmark (Lai et al., 2023), it fixes 44.4%, 14.2%, 20.6%, and 32.1% more data-science bugs than the best prior baseline while cutting prompt tokens by up to 34%, showing that knowledge-graph evidence can outperform unstructured text retrieval in one-shot repair.
@@ 5.4. Analysis-Augmented Generation Enhanced Prompting
Analysis-augmented generation (AAG) feeds a frozen language model with artifacts from program analysis, such as compiler diagnostics, fault localization scores, data flow constraints, failing test logs, or execution traces. These artifacts typically appear in three forms: failing-test evidence that exposes runtime failures, semantics-aware slices that prune unrelated code, and concise execution traces that embed dynamic behavior. Recent systems illustrate each form in turn: D4C (Xu et al., 2025a) injects failing-test evidence, Appatch (Nong et al., 2025) applies semantics-aware slices, and TracePrompt (Haque et al., 2025b) contributes concise execution traces. Each patch is grounded in concrete evidence and is generated in a single pass. Compared to plain zero-shot prompts, AAG consistently improves repair accuracy and keeps edits focused on fault-relevant code, while ensuring that extra context remains within the model window.
D4C aligns the prompt with decoder-only models’ next-token objective and lets GPT-4 complete the whole function while embedding the failing test case and its output (Xu et al., 2025a). Without any fault-localization hints, D4C fixes 180 of 437 single-function bugs in Defects4J, 10% more than the previous state of the art that assumed perfect localization, with only ten samples per bug. The approach relies on dynamic evidence, so its gains diminish when tests give little diagnostic detail.
Appatch applies semantics-aware scoping and adaptive exemplar prompting to vulnerability repair (Nong et al., 2025). It slices the program to include only code relevant to the vulnerability, mines an exemplar database, and selects examples on the fly to guide patch generation. In evaluations on 306 real vulnerabilities, it raised F1 from 44.9% to 57.2% across four LLMs, outpacing baseline prompting by up to 98.9% and surpassing state-of-the-art static approaches by as much as 65.4%. Appatch also fixed seven of eleven zero-day vulnerabilities that none of the baselines could handle. Strengths include reduced context, tailored exemplars, and ensemble validation across models; limitations are the need for a known vulnerability location and extra token overhead from exemplar insertion.
@@
Overview Prompt-based repair has evolved from succinct zero-shot instructions to prompts enriched with exemplar fixes, retrieved project fragments, and static or dynamic analysis outputs. Larger frozen language models consistently repair more bugs without requiring parameter updates, and they become even more effective when supplied with focused auxiliary information, such as highly similar bug-fix pairs, carefully selected repository or API snippets, or compact failing-test logs and execution traces. Conversely, verbose or loosely related context can distract the model and harm accuracy, underscoring the need to balance relevance against prompt length. Most current pipelines still rely on manual wording or heuristic retrieval, but evidence from CEDAR (Nashid et al., 2023) shows that automatic exemplar selection and prompt compression are already practical. Continued research on adaptive artifact gathering, scalable summarization, and systematic evaluation will be essential for making prompt-centric repair a dependable component of continuous-integration workflows and developer tooling.